# To CoT or not to CoT? Chain-of-thought helps mainly on math and symbolic reasoning

### v1

@@ 3.2 Results
4 Results from Experiments
Our analysis of the literature sheds light on the behavior of CoT, but still leaves open questions about the behavior of the newest models and apples-to-apples comparisons across model classes, datasets, and prompting techniques. To further our characterization, we perform a series of experiments on 20 datasets across 14 models in both the zero-shot and few-shot setting to compare performance.
4.1 Experimental Setup
Dataset, Models, Prompts
Table 2 lists the models, datasets, and prompting techniques we consider for our experiments (more details, including the dataset composition of each reasoning category, in Table 4 and Table 5 of Appendix A). We restricted our experiments to English models commonly used and benchmarked on general reasoning datasets. This excludes math-specific language models like DeepSeekMath-Instruct models (Shao et al., 2024) as well as non-English models. We also focus on instruction-tuned language models. Our datasets include those which are widely used in CoT and reasoning literature, including a mix of non-symbolic, semisymbolic, and symbolic reasoning. They span different formats, including multiple-choice, short-answer, and free-response; however, most of these datasets are multiple choice or short answer, as CoT is not typically used in long-form response settings. We also categorize each dataset into a larger category of reasoning required to solve it: Commonsense, Knowledge, Symbolic, Mathematical, and Soft Reasoning. We define Soft Reasoning as questions relying on commonsense and natural language but going beyond simple inferences about these statements. Finally, we explore several prompting strategies for eliciting reasoning from language models, as past work has emphasized the importance of the prompt (Yang et al., 2024). However, we generally found slight performance differences; see Appendix G for details. We therefore focus on prompts similar to Kojima et al. (2022) and Wei et al. (2022) for zero-shot and few-shot settings, respectively, with alterations to improve the model’s ability to produce desired behavior (i.e., formats that allow for easily parsed answers).
Implementation Details
We use a high-throughput inference package, vLLM (Kwon et al., 2023), for the model inference process. We use greedy decoding on all models. Our prompts are taken from the Llama 3.1 evaluations when available (Dubey et al., 2024), and minor adjustments are made to unify prompting strategies. For other datasets, we either use the standard prompt for the dataset from the corresponding original paper or implement our own prompt. [^5] Prompts and outputs for our evaluations are on Huggingface, https://huggingface.co/collections/TAUR-Lab/cot-analysis-project-66bbb9e5e0156e65059895f5 Our answer parser ($\mathrm{extract}$) is tailored to each dataset and model. Specific details about each dataset, its prompts, and answer extractor can be found in Appendix A.
Figure 3: Left: Performance gain from using CoT for each reasoning category. Right: Performance gain from using CoT for each dataset, averaged across models and broken out across 5 representative models. Red lines indicate median improvement. In both plots we see a consistent trend: most improvements from using CoT are from math and symbolic reasoning. This trend remains true across model capabilities.
4.2 Results
Where does zero-shot CoT improve over direct prompts?
By contrast, the mathematical and symbolic categories get larger boosts in improvements alongside symbolic and many semi-symbolic datasets. MATH and GSM8k show gains as large as 41.6% and 66.9%, respectively. The semi-symbolic datasets like ContextHub and MuSR Murder Mysteries show moderate gains. These datasets require the application of logical rules to reach the answer, e.g., first-order logic parsed from simple natural language (ContextHub) or more complex commonsense statements (MuSR Murder Mysteries). All results are shown in the Appendix C.1 as well as a full list of numeric results for both CoT and direct answer prompting in Table LABEL:tab:raw_exp_0_zs_results. We also explored the few-shot setting and found it had little impact on when CoT will help; see Appendix B.
Does the answer format impact where CoT will help?
Many of the commonly-used datasets for problems other than math are multiple choice. We highlight here that CoT has similar performance to direct answer across models for two datasets that are not multiple-choice and contain varying levels of non-symbolic reasoning to answer the question. First, MuSiQue (Trivedi et al., 2022) is a short-form QA task requiring multi-hop reasoning. We consider this a semi-symbolic dataset as the questions have an explicit multi-hop structure. Because answer spans in MuSiQue can be paraphrased in many different ways, we use GPT-4o to judge if two answer spans are equivalent. Despite being semi-symbolic, we see no overall improvement from CoT.
Mostly no, except for MMLU, StrategyQA, and MuSR.
We tested the significance of the improvements from CoT on the 13 datasets in the Knowledge, Soft Reasoning, and Commonsense reasoning categories using paired bootstrapping to assess whether CoT gives a significant improvement. To account for multiple comparisons, we applied a Bonferroni correction, setting the p-value to $0.00027$ to account for the 14 models and 13 datasets. About 38% (58) of the datasets that show a benefit in these three reasoning categories were considered significant. Nearly half of these comparisons (26) are on MMLU and MMLU Pro, which we study more closely in the next section. StrategyQA and MuSR also received a consistent performance boost across 9 and 6 models respectively. StrategyQA is often used to benchmark reasoning methods and is built specifically to get a benefit from methods that decompose the question into steps, so a gain in performance is not unprecedented. MuSR, similarly, was built to have multiple steps of complex natural language reasoning, which may receive benefits from CoT. The remaining datasets that receive significant benefits are spread across the datasets and models.
4.3 Zoom-in: MMLU and MMLU Pro
MMLU MMLU Pro
Breakdown by the presence of equations
We plot the overall CoT delta (performance of CoT minus the performance of direct answer) for both MMLU and MMLU Pro across multiple models between two bins according to this classifier $g$, labeled as “With =” and “Without =”, in Figure 4. We also report the amount of performance gain explained by questions having an “=” vs. not in Appendix E. We find that the majority of the performance gain from CoT on MMLU and MMLU Pro comes from questions that have an “=” in the question or generated responses. Because “=” are usually found in math problems, we equate this to CoT primarily benefiting MMLU and MMLU Pro on the math-related questions with very little to no gain (depending on the model) for non-math questions.
5 Strengths and weaknesses of CoT at formal reasoning
By separating planning and execution in this way, we can test how much a language model can gain from just knowing how to solve a problem (only having a plan), to having a plan and being able to reason about its output (having a plan and solving it with CoT), or to having a plan and then solve it with an external symbolic solver. Given a plan $\mathcal{S}_{\mathrm{plan}}\sim\mathcal{I}^{m}_{\mathrm{planning}}(\mathbf{q})$, we compare the performance of the settings below to evaluate at which stage LM is most effective and falls short.
Figure 5: Prompt variants that separate planning and execution for GSM8K. For all prompt variants besides direct answer and CoT (not shown), we first few-shot prompt an LLM to generate a Python program as a solution plan. For Plan + Direct Solver, the LLM is prompted to directly give an answer from the plan; for Plan + CoT Solver, the LLM is prompted to solve the plan step-by-step with CoT and give an answer; for Plan + Tool Solver, we feed the plan into a Python interpreter.
5.1 Settings Evaluated
Settings 1 and 2: Few-shot direct answer and CoT:
We use the few-shot direct answer and CoT prompts from Section 4.1 as baselines. Figure 5 includes an example of each setting on GSM8K.
Settings 3 and 4: Plan + Direct Solver and Plan + CoT Solver:
Here we use inspiration from Xu et al. (2024a) and generate a symbolic plan using the same strategy as Ye et al. (2023). Specifically, we use a few-shot prompt $\mathcal{I}^{m}_{\mathrm{planning}}$ that is meant to generate a formal specification $\mathcal{S}_{\mathrm{plan}}$ that should be executable by a symbolic solver. [^7] Note that we do not enforce any correctness on the output for this setting. Instead, we assume that given the examples in the prompt, the language model will generate a fairly complete specification. In the same prompt LMs are asked to solve their generated specification $\mathcal{S}_{\mathrm{plan}}$ and derive the final answer $\tilde{\mathbf{y}}\sim p(\mathbf{y}\mid\mathcal{I}_{\mathrm{da}}(\mathcal{S}_{\mathrm{plan}}))$, either directly giving the answer after generating the specification (Plan + Direct Solver) or providing a trace of the plan (step-by-step explanations and tracking of intermediate steps) for the derivation (Plan + CoT Solver). Particularly, $\mathcal{S}_{\mathrm{plan}}$ is a Python program for the math datasets, and is a set of formal specifications in first-order logic for the logical reasoning datasets.
Setting 5: Plan + Tool Solver
We then evaluate how effective CoT can be at performing symbolic computations compared with external symbolic solvers. Following prior work on augmenting LMs with tools for math and logic questions (Ye et al., 2023; Pan et al., 2023; Gao et al., 2023; Chen et al., 2023), we generate $\mathcal{S}_{\mathrm{plan}}$ the same way as in CoT Solver, but now feed in the plan into a symbolic solver (Python interpreter or a SMT Solver), such that $\hat{a}=\mathrm{solve}(\mathcal{S}_{\mathrm{plan}})$.
Evaluation Setup
We compare the performance of each setting on math (GSM8K and GSM8K-Hard) and formal logical reasoning (ContextHub and FOLIO) datasets. Recent LLMs might be overfitting GSM8K by data contamination (Zhang et al., 2024), so we follow Gao et al. (2023) to include GSM8K-Hard, a minimally modified harder version that replaces numbers of GSM8K with larger numbers, to account for the potential contamination.
For Plan + Direct solver and Plan + CoT solver, we use the few-shot prompts from Ye et al. (2023). For Plan + Tool solver, we use state-of-the-art tool-augmented prompting methods using symbolic solvers. Particularly, for GSM8K, we use Program-aided Language Model (Gao et al., 2023, PAL) that executes the LM-generated plan with a Python interpreter. For the logical reasoning dataset, we use Satisfiability-Aided Language Model (Ye et al., 2023, SatLM) that uses automated theorem prover Z3 (De Moura & Bjørner, 2008) to solve the generated specifications. If the LM-generated plan cannot be parsed by the tool, we use random guessing when the question is multiple choice, and mark it incorrect otherwise.
5.2 Evaluation Results
Figure 6 shows the results across a representative selection of models. Detailed numerical results, including the unparseable rates of model-generated plans, can be found in Appendix F.
We argue that these results provide an explanation of why CoT helps on symbolic tasks. While all tasks could feasibly benefit from a detailed description of how to solve each individual question (e.g., a plan in the context of this section), CoT only outperforms direct answer when these steps require a substantial amount of tracing and computation. In these settings, we can see clear performance benefit from using symbolic solvers; CoT appears to be a poor (but universal) approximation to such solvers. When possible, LLMs should be paired with symbolic solvers when solving symbolic tasks to achieve consistently better performance over direct answer and CoT.
6 Discussion and Related Work
Where is CoT helping and why?
Long Horizon Planning
Can we improve CoT further?
Our work treats chain-of-thought variants that explicitly don’t involve multiple inferences. But there is some evidence that using additional calls to LLMs can help (Du et al., 2023; Yao et al., 2023; Besta et al., 2023; Chen et al., 2024). One challenge is that these methods use significantly increased computation; careful benchmarking sometimes reveals that naive techniques are as good as iterative ones (Olausson et al., 2024). However, past theoretical results have shown that Transformers are augmented in a fundamental way by CoT (Liu et al., 2023b; Merrill & Sabharwal, 2024); we believe this does indicate the potential for improved variants of CoT beyond prompt-based CoT. On the other hand, recent methods showing benefit from “internalizing” CoT (Deng et al., 2024) may indicate that explicit generation of intermediate tokens is still not being used to its full potential.
Dataset contamination
7 Conclusion
In this work, we characterize the performance of prompt-based CoT through a meta-analysis of the literature and experiments across different models, datasets, and prompts. We find that CoT predominantly helps on math and formal logic tasks regardless of including examples in the prompt, using different question formats, or running on stronger models. We analyze CoT’s behavior further and find that a majority of the performance gain is consistently attributed to tracing the intermediate steps of a problem, which symbolic solvers are better suited for and thus CoT rarely outperforms them. We believe that CoT remains a powerful technique, but to give improvement across a wider range of NLP tasks, research should move beyond prompt-based CoT to new paradigms like search, interacting agents, or better fine-tuned models.
8 Reproducibility
For our experiments, we provide in-depth details of how we evaluated models on each dataset in Section 4.1 and Appendix A. Furthermore, we release all prompts for every dataset on Huggingface, including per model output and sampling parameters. For our meta-analysis of the literature, we describe our filtering criteria and process of annotating experiments into high-level categories in Section 3 and Appendix D. We also release the full list of papers in our meta-analysis together with extracted experimental comparisons and task category annotations.
Acknowledgments
We acknowledge George Tsoukalas for providing insightful feedback throughout the project. We also thank Kaj Bostrom and Eunsol Choi for reviewing and providing feedback on drafts of the work. This work was partially supported by NSF CAREER Award IIS-2145280 (to Durrett), NSF CAREER Award 2339729 (to Mahowald), the NSF AI Institute for Foundations of Machine Learning (IFML), the Sloan Foundation via a Sloan Research Fellowship, and a grant from Open Philanthropy.
Appendix A Expanded Experimental Details
@@
Appendix B Few-shot Experiments
@@
Appendix C Expanded CoT vs Direct Experimental Results
C.1 Full Zero-shot Results
@@
Claude-3 Haiku 71.8 74.2 73.0 76.1 18.1 48.2 38.6 48.7 55.9 56.6
Claude-3.5 Sonnet 80.7 82.6 83.8 88.8 38.7 59.0 53.2 67.1 67.6 75.7
GPT-4o Mini 78.3 79.9 73.6 83.1 22.9 59.7 48.1 60.9 61.1 63.5
Gemini 1.5 Flash 76.8 73.6 78.2 81.0 27.2 55.7 47.0 59.7 60.6 62.6
Gemini 1.5 Pro 77.3 75.2 80.9 83.8 35.4 58.5 52.9 62.6 64.1 67.8
Gemma 2 9b 71.5 72.9 74.9 76.9 18.5 50.5 46.7 55.8 58.2 60.5
GPT-4o 83.6 84.1 82.9 88.6 36.5 63.3 55.7 68.3 65.9 74.0
Llama 2 7b 50.4 50.1 44.1 46.6 09.3 17.2 22.4 35.4 37.2 37.6
Llama 3.1 70b 81.0 81.6 82.4 85.6 24.9 54.9 49.0 60.0 65.7 69.5
Llama 3.1 8b 70.1 70.9 70.1 74.1 16.0 47.8 34.8 51.6 55.0 56.2
Mistral 7b 62.4 66.0 62.0 64.5 10.9 28.9 41.8 45.0 48.6 49.7
Phi-3 Small 8k 70.2 72.5 76.1 79.7 17.8 47.1 51.2 58.7 57.9 56.4
Qwen 2 72b 79.2 81.7 78.6 84.6 23.9 58.5 48.2 58.7 64.2 65.1
Qwen 2 7b 68.3 70.6 65.2 71.3 15.9 53.5 43.8 52.3 54.4 49.4
Average 73.0 74.0 73.3 77.5 22.6 50.2 45.2 56.1 58.3 60.3
@@
Dataset Type Model zero-shot DA accuracy zero-shot CoT accuracy
@@
SiQA Commonsense Llama 2 7b 49.7 51.1
SiQA Commonsense Mistral 7b 56.9 54.2
SiQA Commonsense Llama 3.1 8b 60.6 59.5
SiQA Commonsense Llama 3.1 70b 63.1 65.0
SiQA Commonsense Gemma 2 9b 58.8 58.7
SiQA Commonsense Phi-3 Small 8k 38.0 37.5
SiQA Commonsense Qwen 2 7b 60.0 61.1
SiQA Commonsense Qwen 2 72b 64.5 61.9
SiQA Commonsense GPT-4o Mini 62.6 62.6
SiQA Commonsense GPT-4o 63.8 63.3
SiQA Commonsense Claude-3 Haiku 60.6 62.0
SiQA Commonsense Claude-3.5 Sonnet 64.0 62.8
SiQA Commonsense Gemini 1.5 Flash 58.5 61.5
SiQA Commonsense Gemini 1.5 Pro 58.5 62.7
@@
C.2 Full Few-shot Results
@@
Dataset Type Model few-shot DA accuracy few-shot CoT accuracy
@@
C.3 Answer Extractor and Average Answer Span Results
@@
Appendix D Quantitative Meta-Analysis
Category Description
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (e.g., PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
Table 17: Categories and their descriptions for the meta-analysis.
Appendix E Performance impacts of “=” on MMLU and MMLU Pro
@@
Appendix F Full results of evaluations on formal reasoning datasets
@@
Appendix G Other CoT prompt variants
G.1 Testing Performance Volatility Across Prompts
Appendix H Example prompts
All prompts can be found on Huggingface at https://huggingface.co/collections/TAUR-Lab/cot-analysis-project-66bbb9e5e0156e65059895f5. We list a few prompt response pairs here in this section.

### v2

@@ Appendix C Expanded Experimental Details
Appendix D Other CoT prompt variants
D.1 Testing Performance Volatility Across Prompts
Appendix E Few-shot Experiments
@@
Appendix F Expanded CoT vs Direct Experimental Results
F.1 Full Zero-shot Results
@@
Claude-3 Haiku 74.3 77.2 73.0 76.1 18.1 48.2 38.6 48.7 55.9 56.6
Claude-3.5 Sonnet 84.3 85.8 83.8 88.8 38.7 59.0 53.2 67.1 67.6 75.7
GPT-4o Mini 81.8 83.2 73.6 83.1 22.9 59.7 48.1 60.9 61.1 63.5
Gemini 1.5 Flash 80.3 76.8 78.2 81.0 27.2 55.7 47.0 59.7 60.6 62.6
Gemini 1.5 Pro 80.4 78.3 80.9 83.8 35.4 58.5 52.9 62.6 64.1 67.8
Gemma 2 9b 75.0 76.1 74.9 76.9 18.5 50.5 46.7 55.8 58.2 60.5
Gpt-4o 87.3 87.7 82.9 88.6 36.5 63.3 55.7 68.3 65.9 74.0
Meta-Llama 2 7b 51.4 50.9 44.1 46.6 9.3 17.2 22.4 35.4 37.2 37.6
Meta-Llama 3.1 70b 84.2 84.7 82.4 85.6 24.9 54.9 49.0 60.0 65.7 69.5
Meta-Llama 3.1 8b 72.9 73.4 70.1 74.1 16.0 47.8 34.8 51.6 55.0 56.2
Mistral 7b 58.3 61.8 62.0 64.5 10.9 28.9 41.8 45.0 48.6 49.7
Phi-3 Small 8k 70.8 72.5 76.1 79.7 17.8 47.1 51.2 58.7 57.9 56.4
Qwen 2 72b 82.9 84.9 78.6 84.6 23.9 58.5 48.2 58.7 64.2 65.1
Qwen 2 7b 64.0 66.1 65.2 71.3 15.9 53.5 43.8 52.3 54.4 49.4
Average 74.8 75.7 73.3 77.5 22.6 50.2 45.2 56.1 58.3 60.3
@@
Dataset Type Model zero-shot CoT accuracy zero-shot DA accuracy
@@
SiQA Commonsense Llama 2 7b 53.4 55.9
SiQA Commonsense Mistral 7b 35.9 33.5
SiQA Commonsense Llama 3.1 8b 73.5 73.5
SiQA Commonsense Llama 3.1 70b 78.7 80.9
SiQA Commonsense Gemma 2 9b 74.9 76.3
SiQA Commonsense Phi-3 Small 8k 38.0 40.4
SiQA Commonsense Qwen 2 7b 37.3 39.3
SiQA Commonsense Qwen 2 72b 80.5 80.4
SiQA Commonsense GPT-4o Mini 79.0 80.0
SiQA Commonsense Gpt-4o 81.9 81.5
SiQA Commonsense Claude-3 Haiku 75.4 74.8
SiQA Commonsense Claude-3.5 Sonnet 79.7 81.0
SiQA Commonsense Gemini 1.5 Flash 74.5 79.1
SiQA Commonsense Gemini 1.5 Pro 73.9 78.2
@@
F.2 Full Few-shot Results
@@
Dataset Type Model few-shot CoT accuracy few-shot DA accuracy
@@
F.3 Answer Extractor and Average Answer Span Results
@@
Appendix G Zoom-in: MMLU and MMLU Pro
MMLU MMLU Pro
Breakdown by the presence of equations
We plot the overall CoT delta (performance of CoT minus the performance of direct answer) for both MMLU and MMLU Pro across multiple models between two bins according to this classifier $g$, labeled as “With =” and “Without =”, in Figure 4. We also report the amount of performance gain explained by questions having an “=” vs. not in Appendix G.1. We find that the majority of the performance gain from CoT on MMLU and MMLU Pro comes from questions that have an “=” in the question or generated responses. Because “=” are usually found in math problems, we equate this to CoT primarily benefiting MMLU and MMLU Pro on the math-related questions with very little to no gain (depending on the model) for non-math questions.
G.1 Performance impacts of “=” on MMLU and MMLU Pro
@@
Appendix H Full results of evaluations on formal reasoning datasets
@@
Appendix I Discussion of Limitations
I.1 Long Horizon Planning
I.2 Dataset contamination
Appendix J Example prompts
We will release all prompts and model responses on our Huggingface repo. We list a few prompt response pairs here in this section.